OpenStudy (anonymous):

A particle moves on the XY plane such that its position is given by t ^{3}-3t and y = t ^{2}-2t for t>= 0, where x and y are in meters and the time t is in seconds. a) Find the magnitude and the direction (0 degree le theta < 360 degree) of the resultant velocity when t = 0.5 . b) Does the particle ever come to a stop? If so, when and where?

OpenStudy (anonymous):

\[ velocity = \sqrt{Vx^2 + Vy^2}\]

OpenStudy (anonymous):

\[lvl=\sqrt{(3(0.5^{2}-3)^{2}+(2(0.5)-2)^{2}}\]

OpenStudy (anonymous):

vx=3t^2-3 vy=2t-2

OpenStudy (anonymous):

\[\tan \theta = \frac{ Vy}{ Vx }\]

OpenStudy (anonymous):

\[lvl=2.4622 \tan−1(-1/-2.25)=23.9624degree\]

OpenStudy (anonymous):

for part B how??

OpenStudy (shubhamsrg):

the particle will come to a stop when magnitude of velocity =0 use this fact for (B)

OpenStudy (anonymous):

any further explanation for part B?

OpenStudy (sirm3d):

the particle comes to a full stop when both horizontal and velocities are equal to zero. set \(v_x=0\) and \(v_y=0\) and solve for \(t\) in each equation. if there is a common value for \(t\), then at that instant, the particle is at rest.

OpenStudy (anonymous):

\[3t ^{2}-3=0 and 2t-2=0?\]

OpenStudy (sirm3d):

thats right

OpenStudy (anonymous):

\[3t ^{2}−3=0\] \[t ^{3}=3t\] \[\frac{ t ^{3} }{ t }=3\]<---- right?

OpenStudy (sirm3d):

nope. \(3t^2-3=0\) \(3t^2=3\) \(t^2=1\) \(t=1\) \(2t-2=0\) \(t=1\) therefore, at \(t=1\) the particle is at rest.

OpenStudy (anonymous):

that all it asking?

OpenStudy (sirm3d):

yes. find the time \(t\) such that the particle is at rest. also, find its position when it is at rest. put \(t=1\) in the equation for \(x\) and \(y\) and you'll get the coordinate of the point where it is at rest.

OpenStudy (anonymous):

x=t^3-3t =1^3-3(1) =1-3 =-2 y=t^2-2t =1^2-2(1) =1-2 =-1

OpenStudy (anonymous):

my answer are - ve not sure if it correct

OpenStudy (sirm3d):

negative values for \(x\) or \(y\) indicate that their position is to the left or below the origin.

OpenStudy (anonymous):

x=-2m and y=-1m ?